Input Variable and NN cuts

Richard St. Denis University of Glasgow/CDF 10 December 2007

H → WW Meeting

[figure]

All Variables

Richard St. Denis, Glasgow University/CDF, 10 Dedember 1

H → WW Meeting

[figure]

LRHWW: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 2

H → WW Meeting

[figure]

LRHWW: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 3

H → WW Meeting

[figure]

LRHWW: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 4

H → WW Meeting

[figure]

LRHWW: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 5

H → WW Meeting

[figure]

LRHWW: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 6

H → WW Meeting

[figure]

LRHWW: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 7

H → WW Meeting

[figure]

LRHWW: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 8

H → WW Meeting

[figure]

LRHWW: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 9

H → WW Meeting

[figure]

dRLeptons: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 10

H → WW Meeting

[figure]

dRLeptons: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 11

H → WW Meeting

[figure]

dRLeptons: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 12

H → WW Meeting

[figure]

dRLeptons: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 13

H → WW Meeting

[figure]

dRLeptons: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 14

H → WW Meeting

[figure]

dRLeptons: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 15

H → WW Meeting

[figure]

dRLeptons: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 16

H → WW Meeting

[figure]

dRLeptons: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 17

H → WW Meeting

[figure]

dPhiLeptons: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 18

H → WW Meeting

[figure]

dPhiLeptons: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 19

H → WW Meeting

[figure]

dPhiLeptons: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 20

H → WW Meeting

[figure]

dPhiLeptons: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 21

H → WW Meeting

[figure]

dPhiLeptons: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 22

H → WW Meeting

[figure]

dPhiLeptons: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 23

H → WW Meeting

[figure]

dPhiLeptons: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 24

H → WW Meeting

[figure]

dPhiLeptons: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 25

H → WW Meeting

[figure]

dimass: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 26

H → WW Meeting

[figure]

dimass: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 27

H → WW Meeting

[figure]

dimass: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 28

H → WW Meeting

[figure]

dimass: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 29

H → WW Meeting

[figure]

dimass: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 30

H → WW Meeting

[figure]

dimass: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 31

H → WW Meeting

[figure]

dimass: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 32

H → WW Meeting

[figure]

dimass: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 33

H → WW Meeting

[figure]

MetSpec: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 34

H → WW Meeting

[figure]

MetSpec: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 35

H → WW Meeting

[figure]

MetSpec: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 36

H → WW Meeting

[figure]

MetSpec: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 37

H → WW Meeting

[figure]

MetSpec: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 38

H → WW Meeting

[figure]

MetSpec: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 39

H → WW Meeting

[figure]

MetSpec: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 40

H → WW Meeting

[figure]

MetSpec: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 41

H → WW Meeting

[figure]

Met: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 42

H → WW Meeting

[figure]

Met: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 43

H → WW Meeting

[figure]

Met: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 44

H → WW Meeting

[figure]

Met: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 45

H → WW Meeting

[figure]

Met: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 46

H → WW Meeting

[figure]

Met: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 47

H → WW Meeting

[figure]

Met: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 48

H → WW Meeting

[figure]

Met: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 49

H → WW Meeting

[figure]

MetSig: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 50

H → WW Meeting

[figure]

MetSig: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 51

H → WW Meeting

[figure]

MetSig: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 52

H → WW Meeting

[figure]

MetSig: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 53

H → WW Meeting

[figure]

MetSig: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 54

H → WW Meeting

[figure]

MetSig: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 55

H → WW Meeting

[figure]

MetSig: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 56

H → WW Meeting

[figure]

MetSig: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 57

H → WW Meeting

[figure]

addEt: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 58

H → WW Meeting

[figure]

addEt: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 59

H → WW Meeting

[figure]

addEt: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 60

H → WW Meeting

[figure]

addEt: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 61

H → WW Meeting

[figure]

addEt: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 62

H → WW Meeting

[figure]

addEt: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 63

H → WW Meeting

[figure]

addEt: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 64

H → WW Meeting

[figure]

addEt: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 65

H → WW Meeting

[figure]

Lep1 E: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 66

H → WW Meeting

[figure]

Lep1 E: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 67

H → WW Meeting

[figure]

Lep1 E: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 68

H → WW Meeting

[figure]

Lep1 E: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 69

H → WW Meeting

[figure]

Lep1 E: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 70

H → WW Meeting

[figure]

Lep1 E: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 71

H → WW Meeting

[figure]

Lep1 E: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 72

H → WW Meeting

[figure]

Lep1 E: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 73

H → WW Meeting

[figure]

MetDelPhi: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 74

H → WW Meeting

[figure]

MetDelPhi: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 75

H → WW Meeting

[figure]

MetDelPhi: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 76

H → WW Meeting

[figure]

MetDelPhi: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 77

H → WW Meeting

[figure]

MetDelPhi: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 78

H → WW Meeting

[figure]

MetDelPhi: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 79

H → WW Meeting

[figure]

MetDelPhi: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 80

H → WW Meeting

[figure]

MetDelPhi: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 81

H → WW Meeting

[figure]

LRZZ: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 82

H → WW Meeting

[figure]

LRZZ: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 83

H → WW Meeting

[figure]

LRZZ: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 84

H → WW Meeting

[figure]

LRZZ: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 85

H → WW Meeting

[figure]

LRZZ: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 86

H → WW Meeting

[figure]

LRZZ: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 87

H → WW Meeting

[figure]

LRZZ: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 88

H → WW Meeting

[figure]

LRZZ: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 89

H → WW Meeting

[figure]

Ht: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 90

H → WW Meeting

[figure]

Ht: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 91

H → WW Meeting

[figure]

Ht: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 92

H → WW Meeting

[figure]

Ht: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 93

H → WW Meeting

[figure]

Ht: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 94

H → WW Meeting

[figure]

Ht: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 95

H → WW Meeting

[figure]

Ht: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 96

H → WW Meeting

[figure]

Ht: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 97

H → WW Meeting

[figure]

SumEt: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 98

H → WW Meeting

[figure]

SumEt: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 99

H → WW Meeting

[figure]

SumEt: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 100

H → WW Meeting

[figure]

SumEt: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 101

H → WW Meeting

[figure]

SumEt: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 102

H → WW Meeting

[figure]

SumEt: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 103

H → WW Meeting

[figure]

SumEt: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 104

H → WW Meeting

[figure]

SumEt: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 105

H → WW Meeting

[figure]

Lep2 E: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 106

H → WW Meeting

[figure]

Lep2 E: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 107

H → WW Meeting

[figure]

Lep2 E: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 108

H → WW Meeting

[figure]

Lep2 E: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 109

H → WW Meeting

[figure]

Lep2 E: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 110

H → WW Meeting

[figure]

Lep2 E: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 111

H → WW Meeting

[figure]

Lep2 E: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 112

H → WW Meeting

[figure]

Lep2 E: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 113

H → WW Meeting

[figure]

LRWW: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 114

H → WW Meeting

[figure]

LRWW: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 115

H → WW Meeting

[figure]

LRWW: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 116

H → WW Meeting

[figure]

LRWW: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 117

H → WW Meeting

[figure]

LRWW: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 118

H → WW Meeting

[figure]

LRWW: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 119

H → WW Meeting

[figure]

LRWW: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 120

H → WW Meeting

[figure]

LRWW: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 121

H → WW Meeting

[figure]

LRWj: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 122

H → WW Meeting

[figure]

LRWj: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 123

H → WW Meeting

[figure]

LRWj: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 124

H → WW Meeting

[figure]

LRWj: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 125

H → WW Meeting

[figure]

LRWj: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 126

H → WW Meeting

[figure]

LRWj: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 127

H → WW Meeting

[figure]

LRWj: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 128

H → WW Meeting

[figure]

LRWj: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 129

H → WW Meeting

[figure]

LRWg: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 130

H → WW Meeting

[figure]

LRWg: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 131

H → WW Meeting

[figure]

LRWg: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 132

H → WW Meeting

[figure]

LRWg: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 133

H → WW Meeting

[figure]

LRWg: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 134

H → WW Meeting

[figure]

LRWg: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 135

H → WW Meeting

[figure]

LRWg: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 136

H → WW Meeting

[figure]

LRWg: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 137

H → WW Meeting

[figure]

Jet1 Et: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 138

H → WW Meeting

[figure]

Jet1 Et: High S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 139

H → WW Meeting

[figure]

Jet1 Et: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 140

H → WW Meeting

[figure]

Jet1 Et: High S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 141

H → WW Meeting

[figure]

Jet1 Et: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 142

H → WW Meeting

[figure]

Jet1 Et: Low S/B 0 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 143

H → WW Meeting

[figure]

Jet1 Et: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 144

H → WW Meeting

[figure]

Jet1 Et: Low S/B 1 Jet

Richard St. Denis, Glasgow University/CDF, 10 Dedember 145